# Nonreciprocal quantum phase transition in a spinning microwave magnonic system

### v1

@@ II The model
with $\Pi_{\pm}=\pm\tilde{\Delta}_{a}-G$, $\Theta_{\pm}=-\gamma\pm K$Im$\left[M^{2}\right]$ and $\Xi_{\pm}=\pm\tilde{\Delta}_{m}-K$Re$\left[M^{2}\right]$. For a given solution of $M$ in Eq. (5), only if all eigenvalues of the matrix $\mathbf{U}$ have negative real parts, the solution is said to be stable, otherwise this solution is unstable. After numerically carrying out this stability analysis [see Fig. 2(a)-2(c) and related discussions], we find that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter space, while the solutions $\left|M\right|_{-}^{2}$ and $\left|M\right|_{+}^{2}$ are stable in some parameter space, which suggests that we only need to apply the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ to investigate the steady-state quantum phase transitions. In addition, to research the variances of magnon mode quadratures, we define the time-dependent covariance matrix $V\left(t\right)$ with $V_{ij}\left(t\right)=\left\langle f_{i}\left(t\right)f_{j}\left(t^{\prime}\right)+f_{j}\left(t^{\prime}\right)f_{i}\left(t\right)\right\rangle/2$ $\left(i,j=1,2,3,4\right)$. From Eq. (11), one can easily find the solutions of $V\left(\infty\right)$ by solving the so-called Lyapunov equation $\mathbf{UV}+\mathbf{VU}^{T}=-\mathbf{D}$ with $\mathbf{D}$ being the diffusion matrix, defined as $\mathbf{D}_{ij}\left(t-t^{\prime}\right)=\left\langle\delta\mathbf{O}_{in,i}\left(t\right)\delta\mathbf{O}_{in,j}\left(t^{\prime}\right)+\delta\mathbf{O}_{in,j}\left(t^{\prime}\right)\delta\mathbf{O}_{in,i}\left(t\right)\right\rangle/2$. The diagonal elements of the matrix $\mathbf{V}$ associate with the variance of quadratures. In our model, the variance of magnon mode quadrature is given by $\left\langle\delta m^{\dagger}\delta m\right\rangle=\left[\left(\mathbf{V}_{33}+\mathbf{V}_{44}\right)-1\right]/2$.
@@ III Nonreciprocal quantum phase transition
In this section, we explore how to realize an nonreciprocal QPT via the spinning of the resonator. We begin by employing the standard stability analysis to show the phase diagrams against the reduced driving strength $G/\kappa$ and the ratio $\Delta_{m}/\tilde{\Delta}_{c}$. As illustrated in Figs. 2(a)-2(c), each steady-state phase diagram exists three different regions corresponding to different phases, i.e., PSP, PSBP, and BP. The vertical black solid line ($G=G_{c1}$) is the boundary between PSP and BP in the case of $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=Gc1}$, and then the blue and red solid curves ($G=G_{c2}$) are the boundaries between PSP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\geq\beta|_{G=Gc1}$, BP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=Gc1}$, respectively. These manifest that the boundaries between different phases are determined by the critical driving strengths $G_{c1}$ and $G_{c2}$. In particular, the intersection point of three boundaries emerges when $G_{c1}=G_{c2}$ at $\Delta_{m}/\tilde{\Delta}_{a}=\beta|_{G=Gc1}$. More interestingly, we find that different spinning directions (the left or right) give rise to QPT at different critical driving strength. Compared with the stationary case (i.e., no spinning with $\Delta_{F}=0$), the areas of three phases have a remarkable change for $\Delta_{F}>0$ and $\Delta_{F}<0$. Therefore, this QPT are highly controllable and can be tuned by the driving direction (or the spinning direction) and the spinning speed of the resonator. As a result, the nonreciprocal QPT can be achieved in this system.
@@
In Figs. 3(a)-2(c), we further display the dynamical behaviors of the scaled mean magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ to reveal the macroscopic magnon excitations for different phases. Here the parameter values for each phase have marked with different-colored dots in Fig. 2(b). When the system is in BP [see Fig. 3(a)], we find that both the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ are stable and the steady-state magnon occupation can be dominated by the initial condition of the system. This situation indicates that the parity symmetry in BP can be conserved or broken with appropriate initial conditions. It is noticed that PSP and BP will cannot be distinguished by only tracking the dynamics of the system with an initial state near zero since the solution $\left|M\right|_{0}^{2}$ is stable for both phases. In this case, the initial state should be set far away from zero to distinguish the two phases and observe the phase transition. From Fig. 3(b), we see that only the nontrivial solution $\left|M\right|_{+}^{2}$ becomes stable in PSBP. The system has the macroscopic magnon excitation in this phase and thus is parity-symmetry broken. As described in Fig. 3(c), there is no macroscopic magnon excitation in PSP, that is to say, only the solution $\left|M\right|_{0}^{2}$ is stable and then the system is parity symmetric. Besides, an interesting feature is that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter range since the four eigenvalues of the matrix $\mathbf{U}$ in Eq. (12) have at least one nonnegative real part.
@@
Figure 4: Normalized steady-state magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ is plotted as the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3$, 00, and $0.3$. The solid curves correspond to the analytical results in Eq. (6) and the circles correspond to the numerical results obtained by Eq. (LABEL:7). (a) $\Delta_{m}/\kappa=4$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=0.2+0.2i$, (b) $\Delta_{m}/\kappa=2.2$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=10+10i$. The other parameters are $\Delta_{c}/\kappa=3$, $J/\kappa=2.5$, and $\gamma/\kappa=1.$
Next, we reveal more characteristics of the nonreciprocal QPT through numerically observing the behavior of the correlated fluctuation of the magnon operators $\delta m^{\dagger}$ and $\delta m$. As shown in Fig. 5, we plot the expectation value $\left\langle\delta m^{\dagger}\delta m\right\rangle$ of correlated fluctuation as a function of the scaled coupling strength $G/\kappa$ for different values of $\Delta_{F}$. There is a common characteristic in both Figures 5(a) and (5b), that is, the mean correlated fluctuation $\left\langle\delta m^{\dagger}\delta m\right\rangle$ is close to zero when the driving strength $G$ is away from the critical value, but Figure 5(a) (Figure 5(b)) is of a asymptotic (catastrophic) divergent behavior around the critical point, which is connected with the second-order QPT from PSP to PSBP (the first-order QPT from PSP to BP). From Fig. 5, we notice especially that spinning the resonator decreases the critical driving strength for $\Delta_{F}<0$ or increases it for $\Delta_{F}>0$, compared with the stationary-resonator case ($\Delta_{F}=0$). Besides, associating Fig. 4 with Fig. 5 finds that the the critical positions where the QPT occurs in Fig. 5(a) (Fig. 5(b)) completely agree with those in Fig. 4(a) (Fig. 4(b)) as we expected. The results give further evidence that the nonreciprocal QPT can be realized in the model.
Figure 5: Mean correlated fluctuation $\left\langle\delta m^{\dagger}\delta m\right\rangle$ as a function of the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3,0,$and $0.3$. The other parameters are the same with Figure 4.
Finally, to quantitatively describe the effect of spinning on the nonreciprocal QPT, we introduce the isolation parameter
@@
=$\mathfrak{R}=\left\{\begin{array}[]{cc}0,\text{ }&\left|M\right|^{2}\left(\Delta_{F}\right)=0\\ \left|\frac{\left|M\right|^{2}\left(\Delta_{F}<0\right)-\left|M\right|^{2}\left(\Delta_{F}>0\right)}{\left|M\right|^{2}\left(\Delta_{F}<0\right)+\left|M\right|^{2}\left(\Delta_{F}>0\right)}\right|.&The\text{ }other\end{array}\right.$ (13)
@@
For a QPT without the spinning resonator (i.e., a conventional reciprocal QPT), the isolation parameter is $\mathfrak{R}=0$. A nonzero $\mathfrak{R}$ denotes the emergence of nonreciprocity in the phase transition. The higher the isolation parameter $\mathfrak{R}$ is, the stronger the nonreciprocity of the QPT is. Especially, $\mathfrak{R}=1$ corresponds to an ideal nonreciprocal QPT. The isolation parameter $\mathfrak{R}$ as the function of the Fizeau shift $\left|\Delta_{F}\right|/\kappa$ is shown Figs. 6(a) and 6(b) for the second-order and first-order phase transitions, respectively. As expected, the isolation parameter is $\mathfrak{R}=0$ for the QPT with a stationary resonator (i.e., $\Delta_{F}=0$). While it is clear that, not only the nonreciprocal QPT occurs in a remarkably broad parameter range, but also the region of the ideal nonreciprocal QPT enlarges with the increasing of the Fizeau shift $\left|\Delta_{F}\right|$. Further, the ideal nonreciprocal QPT exists obviously two boundaries, one boundary condition (the left black curve) satisfies $G=G_{c2}\left(\Delta_{F}<0\right)$ and the other one (the right black curve) is $G=G_{c2}\left(\Delta_{F}>0\right)$ in Fig. 5(a). In the same way, there exists a similar characteristic in Fig. 5(b). An interesting feature in Fig. 6 is that in the vicinity of the left black curses, the isolation parameter $\mathfrak{R}$ experiences a sudden change from zero to one. That is because both mean magnon numbers $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are zero when the driving strength is less than the critical strength, However, when the driving strength is more than the the critical strength, $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is still zero. In this situation, we see that the isolation parameter $\mathfrak{R}$ undergoes a gradually change from one to zero. It can be well understood that $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is zero when the driving strength is less than the critical strength, on the other hand, when the driving strength is more than the critical strength, both $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are not only nonzero, but also gradually increasing with the driving strength increasing.
@@
Figure 6: Isolation parameter $\mathfrak{R}$ versus the scaled Fizeau shift $\left|\Delta_{F}\right|/\kappa$ and normalized driving strength $G/\kappa$. The left and right black curves correspond to the boundary conditions in Eqs. (8) and (9), respectively. The parameters are the same as those in Figure 2.
IV Conclusions

### v2

@@ II The model
with $\Pi_{\pm}=\pm\tilde{\Delta}_{a}-G$, $\Theta_{\pm}=-\gamma\pm K$Im$\left[M^{2}\right]$ and $\Xi_{\pm}=\pm\tilde{\Delta}_{m}-K$Re$\left[M^{2}\right]$. For a given solution of $M$ in Eq. (5), only if all eigenvalues of the matrix $\mathbf{U}$ have negative real parts, the solution is said to be stable, otherwise this solution is unstable 65 . After numerically carrying out this stability analysis [see Fig. 2 and related discussions], we find that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter space, while the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ are stable in some parameter space, which suggests that we only need to apply the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ to investigate the steady-state quantum phase transition. In addition, to research the variances of magnon mode quadratures, we define the time-dependent covariance matrix $\mathbf{V}\left(t\right)$ with $V_{\text{ij}}\left(t\right)=\left\langle f_{i}\left(t\right)f_{j}\left(t^{\prime}\right)+f_{j}\left(t^{\prime}\right)f_{i}\left(t\right)\right\rangle/2$ $\left(i,j=1,2,3,4\right)$. From Eq. (11), one can easily find the solutions of $V\left(\infty\right)$ by solving the so-called Lyapunov equation $\mathbf{UV}+\mathbf{VU}^{T}=-\mathbf{D}$ with $\mathbf{D}$ being the diffusion matrix, defined as $D_{\text{ij}}\left(t-t^{\prime}\right)=\left\langle\delta O_{\text{in},i}\left(t\right)\delta O_{\text{in},j}\left(t^{\prime}\right)+\delta O_{\text{in},j}\left(t^{\prime}\right)\delta O_{\text{in},i}\left(t\right)\right\rangle/2$. The diagonal elements of the matrix $\mathbf{V}$ associate with the variance of quadratures. In our model, the variance of magnon mode quadrature is given by $\langle\delta m^{\dagger}\delta m\rangle$=$\left[\left(V_{33}+V_{44}\right)-1\right]/2$ 18 .
@@ III Nonreciprocal quantum phase transition
In this section, we explore how to realize a nonreciprocal QPT via the spinning of the resonator. We begin by employing the standard stability analysis to show the phase diagrams against the reduced driving strength $G/\kappa$ and the ratio $\Delta_{m}/\tilde{\Delta}_{c}$. As illustrated in Figs. 2(a)-2(c), each steady-state phase diagram exists three different regions corresponding to different phases, i.e., PSP, PSBP, and BP. It is worth noting that the closed system (i.e., excluding system dissipations) without drive (for example, standard Dicke model 2 ) can experience an equilibrium QPT, whose occurrence is often accompanied by a mutation of the ground state of the system. Different from the equilibrium QPT, the nonequilibrium QPT occurs in the driven-dissipative system, heralding the alteration of the steady state (rather than the ground state) of the system. In this work, the cavity-magnon system is driven-dissipative, and each point in the phase diagrams corresponds to stable states of the system [cf. Figs. 2(a)-2(c)]. In the PSP (PSBP), the system is in the steady state without (with) macroscopic magnon excitation. However, the system has two steady states in the BP, one without macroscopic magnon excitation and the other with macroscopic magnon excitation (cf. Fig. 3 and related discussions). In Fig. 2, the vertical black solid line ($G=G_{\text{c1}}$) is the boundary between PSP and BP in the case of $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=G_{\text{c1}}}$, and then the blue and red solid curves ($G=G_{\text{c2}}$) are the boundaries between PSP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\geq\beta|_{G=G_{\text{c1}}}$, BP and PSBP for $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta|_{G=G_{\text{c1}}}$, respectively. These manifest that the boundaries between different phases are determined by the critical driving strengths $G_{\text{c1}}$ and $G_{\text{c2}}$. In particular, the intersection point of three boundaries emerges when $G_{\text{c1}}=G_{\text{c2}}$ at $\Delta_{m}/\tilde{\Delta}_{a}=\beta|_{G=G_{\text{c1}}}$. More interestingly, we find that different spinning directions (the left or right) give rise to QPT at different critical driving strength. Compared with the stationary case (i.e., no spinning with $\Delta_{F}=0$), the areas of three phases have a remarkable change for $\Delta_{F}>0$ and $\Delta_{F}<0$. Therefore, this QPT are highly controllable and can be tuned by the driving direction (or the spinning direction) and the spinning speed of the resonator. As a result, the nonreciprocal QPT can be achieved in this system.
@@
In Figs. 3(a)-3(c), we further display the dynamical behaviors of the scaled mean magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ to reveal the macroscopic magnon excitations for different phases. Here the parameter values for each phase have marked with different-colored dots in Fig. 2(b). When the system is in BP [see Fig. 3(a)], we find that both the solutions $\left|M\right|_{0}^{2}$ and $\left|M\right|_{+}^{2}$ are stable and the steady-state magnon occupation can be dominated by the initial condition of the system. When the system stabilizes to the steady state $\left|M\right|_{0}^{2}$ $(=0)$, the steady state has the even parity, i.e., the parity symmetry of the system is conserved. Conversely, since $\left|M\right|_{+}^{2}$ can be odd or even (corresponding to the odd or even parity), it is difficult to say that the parity of the steady state is odd or even in this case of $\left|M\right|_{+}^{2}$. For the steady state with $\left|M\right|_{+}^{2}$, the parity symmetry is broken. It is noticed that PSP and BP cannot be distinguished by only tracking the dynamics of the system with an initial state near zero since the solution $\left|M\right|_{0}^{2}$ is stable for both phases. In this case, the initial state should be set far away from zero to distinguish the two phases and observe the phase transition. From Fig. 3(b), we see that only the nontrivial solution $\left|M\right|_{+}^{2}$ is stable in PSBP, so that the macroscopic magnon excitation (i.e., $\left|M\right|_{+}^{2}$) can continuously varies and the parity symmetry of the system is broken in this situation. As described in Fig. 3(c), there is no macroscopic magnon excitation in PSP, that is to say, only the solution $\left|M\right|_{0}^{2}$ $(=0)$ is stable and then the system is parity-symmetric. Besides, an interesting feature is that the nontrivial solution $\left|M\right|_{-}^{2}$ is always unstable in the whole parameter range since the four eigenvalues of the matrix $\mathbf{U}$ in Eq. (12) have at least one nonnegative real part.
@@
Figure 4: Normalized steady-state magnon number $\left|M\right|^{2}/\left(\gamma/K\right)$ is plotted as the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3$, 00, and $0.3$. The solid curves correspond to the analytical results in Eq. (6) and the circles correspond to the numerical results obtained by Eq. (5). (a) $\Delta_{m}/\kappa=4$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=0.2+0.2i$, (b) $\Delta_{m}/\kappa=2.2$ and initial conditions $\left\langle a\right\rangle_{t=0}/\sqrt{\gamma/K}=\left\langle b\right\rangle_{t=0}/\sqrt{\gamma/K}=10+10i$. The other parameters are $\Delta_{c}/\kappa=3$, $J/\kappa=2.5$, and $\gamma/\kappa=1.$
Next, to quantitatively describe the nonreciprocal QPT, we introduce the isolation parameter 47a
@@
=$\mathfrak{R}=\left\{\begin{array}[]{cc}0,\text{ }&\left|M\right|^{2}\left(\Delta_{F}>0\right){\small=}\left|M\right|^{2}\left(\Delta_{F}<0\right),\\ \left|\frac{\left|M\right|^{2}\left(\Delta_{F}<0\right)-\left|M\right|^{2}\left(\Delta_{F}>0\right)}{\left|M\right|^{2}\left(\Delta_{F}<0\right)+\left|M\right|^{2}\left(\Delta_{F}>0\right)}\right|,&\left|M\right|^{2}\left(\Delta_{F}>0\right){\small\neq}\left|M\right|^{2}\left(\Delta_{F}<0\right).\end{array}\right.$ (13)
@@
For a QPT without the spinning resonator (i.e., a conventional reciprocal QPT), the isolation parameter is $\mathfrak{R}=0$. A nonzero $\mathfrak{R}$ denotes the emergence of nonreciprocity in the phase transition. The higher the isolation parameter $\mathfrak{R}$ is, the stronger the nonreciprocity of the QPT is. Especially, $\mathfrak{R}=1$ corresponds to an ideal nonreciprocal QPT. The isolation parameter $\mathfrak{R}$ as the function of the Fizeau shift $\left|\Delta_{F}\right|/\kappa$ is shown in Figs. 5(a) and 5(b) for the second-order and first-order phase transitions, respectively. As expected, the isolation parameter is $\mathfrak{R}=0$ for the QPT with a stationary resonator (i.e., $\Delta_{F}=0$). While it is clear that, not only the nonreciprocal QPT occurs in a remarkably broad parameter range, but also the region of the ideal nonreciprocal QPT enlarges with the increasing of the Fizeau shift $\left|\Delta_{F}\right|$. Further, the ideal nonreciprocal QPT exists obviously two boundaries, one boundary condition (the left black curve) satisfies $G=G_{\text{c2}}\left(\Delta_{F}<0\right)$ and the other one (the right black curve) satisfies $G=G_{\text{c2}}\left(\Delta_{F}>0\right)$ in Fig. 5(a). In the same way, there exists a similar characteristic in Fig. 5(b), where the boundary condition (the left black curve) obeys $G=G_{\text{c1}}\left(\Delta_{F}<0\right)$ and the other one (the right black curve) obeys $G=G_{\text{c1}}\left(\Delta_{F}>0\right)$. An interesting feature in Fig. 5 is that in the vicinity of the left black curses, the isolation parameter $\mathfrak{R}$ experiences a sudden change from zero to one. This is because both mean magnon numbers $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are zero when the driving strength is less than the critical strength. However, when the driving strength is more than the critical strength, $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is still zero. In this situation, we see that the isolation parameter $\mathfrak{R}$ undergoes a gradually change from one to zero. It can be well understood that $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ is nonzero and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ is zero when the driving strength is less than the critical strength. On the other hand, when the driving strength is more than the critical strength, both $\left|M\right|^{2}\left(\Delta_{F}<0\right)$ and $\left|M\right|^{2}\left(\Delta_{F}>0\right)$ are not only nonzero, but also gradually increasing with the driving strength increasing.
@@
Figure 5: Isolation parameter $\mathfrak{R}$ versus the scaled Fizeau shift $\left|\Delta_{F}\right|/\kappa$ and normalized driving strength $G/\kappa$. The left and right black curves correspond to the boundary conditions in Eqs. (8) and (9), respectively. The parameters are the same as those in Fig. 2.
Finally, we reveal more characteristics of the nonreciprocal QPT through numerically simulating the behavior of the correlated fluctuations of the magnon operators $\delta m^{\dagger}$ and $\delta m$. As shown in Fig. 6, we plot the expectation value $\langle\delta m^{\dagger}\delta m\rangle$ of correlated fluctuation as a function of the scaled coupling strength $G/\kappa$ for different values of $\Delta_{F}$. There is a common characteristic in both Figs. 6(a) and 6(b), that is, the mean correlated fluctuation $\langle\delta m^{\dagger}\delta m\rangle$ is close to zero when the driving strength $G$ is away from the critical value, but Fig. 6(a) (Fig. 6(b)) is of a asymptotic (catastrophic) divergent behavior around the critical point, which is connected with the second-order QPT from PSP to PSBP (the first-order QPT from PSP to BP). From Fig. 6, we notice especially that spinning the resonator decreases the critical driving strength for $\Delta_{F}<0$ or increases it for $\Delta_{F}>0$, compared with the stationary-resonator case ($\Delta_{F}=0$). Besides, we find from Figs. 4 and 6 that the critical positions where the QPT occurs in Fig. 6(a) (Fig. 6(b)) completely agree with those in Fig. 4(a) (Fig. 4(b)) as we expected. The results give further evidences for the nonreciprocal QPT.
Figure 6: Mean correlated fluctuation $\langle\delta m^{\dagger}\delta m\rangle$ as a function of the scaled driving strength $G/\kappa$ for three different Fizeau shifts $\Delta_{F}/\kappa=-0.3,0,$and $0.3$. The other parameters are the same with Fig. 4.
IV Discussions and conclusions
Before concluding, we briefly discuss the experimental feasibility of the present proposal. In cavity magnonics, the decay rate $\kappa$ of the microwave cavity, the damping rate $\gamma$ of the magnon mode, as well as the coupling strength $J$ between the photon and magnon modes are of the order $1$ MHz (e.g., $\kappa/2\pi=2.04$ MHz, $\gamma/2\pi=1.49$ MHz, and $J/2\pi=8.17$ MHz in Ref. 67 ). In addition, the nonlinear coefficient $K$ of the magnon Kerr effect is of the order $1$ nHz 03a , and the thermodynamic limit (i.e., the weak nonlinearity limit) $\gamma/K\rightarrow\infty$ can be easily satisfied. By controlling the strength of the flux drive on a Josephson parametric amplifier, the amplitude $G$ of the parametric driven can be adjusted from 00 to $6$ MHz 68 . As for the nonreciprocity induced by the Fizeau shift, it is proportional to the spinning speed of the resonator, for instance, a spinning speed of $12.9$ kHz can bring about the Fizeau shift $\Delta_{F}$ $\approx 1.14$ MHz 69 . Therefore, our proposal is likely to be achieved in in cavity-magnon systems with current technologies.